OpenStudy (anonymous):

Five different written driving tests are administered by the Motor Vehicle Department. One of these five tests is selected at random for each applicant for a driver's license. A group consisting of two women and three men apply for a license. (Round your answers to three decimal places.) (a) What is the probability that exactly two of the five will take the same test?

OpenStudy (ybarrap):

P = P(taking a particular test) = 1/5 P(2 people take same test out of 2) = 1/5 * 1/5 P( 3 people take different tests given first two take same test) = (4/5) * (3/5) * (2/5) Ways that 2 people out of 5 can take the same test = 5C2 = 5!/(2!3!) P(2 take same test and 3 take different tests) = 5C2 (1/5)^2 * (4/5) *(3/5)*(2/5) = 48/625 = .0768 Ans .0768